OpenStudy (anonymous):

The number of bricks a bricklayer can lay, b, varies directly with the amount of time the bricklayer is laying bricks, t. George can lay 45 bricks in 30 min. Which direct linear variation equation models the information? A. b = 0.3t B. b = 0.6t C. b = 1.5t D. b = 3t

OpenStudy (anonymous):

2 medals

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

can u help me

OpenStudy (anonymous):

Let k be bricks per minute. Solve the following for k. 45 = k 30 k = 3/2 b = (3/2) t or 1.5 t

OpenStudy (anonymous):

t=2b/3